Grade 12th passIIT JEE Entrance Exam

Solve this x(7x-5)=12?

TEJKIRAT SINGH
11 Years agoGrade 12th pass

1 Answer

Vikas TU
ApprovedApproved Tutor Answer11 Years ago

7x^2 – 5x = 12

7x^2 – 5x - 12=0

7x^2 – 12x +7x – 12 = 0
x(7x-12) + (7x-12) = 0
(x+1)(7x-12)=0